OpenStudy (anonymous):

Is (8, -1) a solution of the system, 2x + 3y = 13 and x - y = 9?

OpenStudy (anonymous):

Do you know how to check if a coordinate is a solution of a system?

OpenStudy (austinl):

You have (8, -1) --> (x, y) So what you can do is plug in the values for the equations that you have, if it holds true in both instances, then it is a solution. 2x+3y=13 2(8)+3(-1)=13 16-3=13 13=13 TRUE x-y=9 8-(-1)=9 8+1=9 9=9 TRUE So therefore, it is a solution! :D